Grade 12Algebra

Solve the following:
Laxmi’s father is 49 years old. He is 4 years older than three times Laxmi’s age. What is Laxmi's age?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

LetLaxmi’s age to be= y years old
From the question it is given that,
Lakshmi’s father is4 years older than three times of her age
= 3 ×Laxmi’s age + 4 = Age of Lakshmi’s father= (3 × y) + 4 = 49
= 3y + 4 = 49
By transposing 4 from LHS to RHS it becomes -4 = 3y = 49 - 4
= 3y = 45
Divide both the side by 3
= 3y/3 = 45/3
= y = 15
So, Lakshmi’s age is 15 years

Thanks